OpenStudy (anonymous):

An 18-foot pole is leaning against a wall. It touches the wall 16 feet above the ground. Find the measure of the angle that the pole makes with the ground. Round your answer to the nearest degree.

OpenStudy (bahrom7893):



OpenStudy (bahrom7893):

Sin(theta) = 16/18 = 8/9

OpenStudy (bahrom7893):

theta = ArcSin(8/9)